LASER INDUCED FLUORESCENCE SPECTROSCOPY OF THE SiNSi RADICAL I: THE C̃ ^{2}Δ_{u} – X̃ ^{2}Π_{g} TRANSITION

\underline{MASARU FUKUSHIMA}, TAKASHI ISHIWATA, Faculty of Information Sciences, Hiroshima City University, Hiroshima 731-3194, Japan; CHIHAYA MOTOYOSHI, YOSHIHIRO SUMIYOSHI, and YASUKI ENDO, Department of Basic Science, The University of Tokyo, Meguro 153-8902, Japan.

We have generated SiNSi in a supersonic free jet expansion, and measured the laser induced fluorescence ( LIF ) spectrum in the ultraviolet ( UV ) region. Dispersed fluorescence ( DF ) spectra from the single vibronic levels ( SVL ) have also been recorded by exciting each observed vibronic band. On the basis of the rotational analyses of the observed vibronic bands and the appearance the vibrational structure in the DF spectra, several bands were attributed to the C̃ ^{2}Δ_{u} – X̃ 2Π_{g} system of SiNSi. From the analysis of the 0_{0}^{0} band with an origin of 29,261.61 cm^{−1}, the effective spin-orbit constants were determined to be 140.59 and 2.623 cm^{−1} for the X̃ ^{2}Π_{g} and C̃ ^{2}Δ_{u} states, respectively. In total, ten vibronic levels including the zero vibrational level were assigned to the C̃ ^{2}Δ_{u} state, and the vibonic structure was analyzed considering the Renner-Teller interaction.